OpenStudy (anonymous):

0.5t+0.25(t+16)=4+0.75t

OpenStudy (anonymous):

u forgot the t with the 50

OpenStudy (anonymous):

ok so its 0.5t+0.25t+4=4+0.75t 50t+25t+400=400+75t 75t-75t=0 0=0 no solution

OpenStudy (anonymous):

HOW IS IT NO SOLUTION WHEN IT =0 sorry caps

OpenStudy (anonymous):

lol its ok oh because the t canceled out to 0 so that means that its no solution

OpenStudy (anonymous):

my teacher said it would be identity because it = to something

OpenStudy (anonymous):

yes its identity property because 0=0 but also its no solution

OpenStudy (anonymous):

ok XD

OpenStudy (anonymous):

yes i hope im right!! LOL